# Topological Optimal Transport for Geometric Cycle Matching

### v1

@@ 2.2 Optimal transport on persistence diagrams
where each $\delta_{x_{i}}$ denotes a Dirac supported at $x_{i}=(b_{i},d_{i})$.
Let $\overline{\Lambda}:=\Lambda\cup\{\partial_{\Lambda}\}$ be obtained by adjoining to $\Lambda$ the virtual point $\partial_{\Lambda}$, with the resulting space endowed with the disjoint union topology. Given two measure persistence diagrams $\mu$ and $\mu^{\prime}$, we say a Radon measure $\pi$ on $\overline{\Lambda}\times\overline{\Lambda}$ is admissible if it has marginals $\mu$ and $\mu^{\prime}$ (we consider $\mu$ and $\mu^{\prime}$ to be measures on $\overline{\Lambda}$ which are supported on $\Lambda$) and it satisfies the additional constraint $\pi((\partial_{\Lambda},\partial_{\Lambda}))=0$, and write $\pi\in\Pi_{\mathrm{adm}}(\mu,\mu^{\prime})$. Then for $p\in[1,\infty)$, we define the $p$-Wasserstein distance between measure persistence diagrams [17] as:
$d_{\mathrm{W},p}^{\mathrm{MPD}}(\mu,\mu^{\prime})=\inf_{\pi\in\Pi_{\rm adm}(\mu,\mu^{\prime})}\left(\int_{\overline{\Lambda}\times\overline{\Lambda}}\|x-x^{\prime}\|^{p}_{p}\>\mathrm{d}\pi(x,x^{\prime})\right)^{1/p}.$ (4)
If $\nu_{D}$ and $\nu_{D^{\prime}}$ arise from finite persistence diagrams $D$ and $D^{\prime}$, respectively, as in (3), then it is not hard to show that $d_{\mathrm{W},p}^{PD}(D,D^{\prime})=d_{\mathrm{W},p}^{\mathrm{MPD}}(\nu_{D},\nu_{D^{\prime}})$ [17, Prop. 3.2]. In general, $d_{\mathrm{W},p}^{\mathrm{MPD}}$ is finite, provided that we work in the space of measure persistence diagrams $\mu$ with finite $p$-persistence, in the sense that
[…]
In what follows, for any given $p$, we will restrict to the space of measure persistence diagrams $\mu$ with finite $p$-persistence. With an abuse of notation, we will still indicate their space as MPD.
@@ 2.3 Gromov-Wasserstein and co-optimal transport distances
A metric measure space (mm-space) [43, 32] is a triple $M=(X,d,\mu)$, where $X$ is a space endowed with a complete separable metric $d$ and with a fully supported Borel measure $\mu$. Given two mm-spaces $M=(X,d,\mu)$ and $M^{\prime}=(X^{\prime},d^{\prime},\mu^{\prime})$ and a coupling $\pi\in\Pi(\mu,\mu^{\prime})$ and $p\in[1,\infty)$, we introduce the $p$-distortion $\mathrm{dis}_{p}^{\mathrm{GW}}(\pi)$:
@@
For our purposes, we relax the requirement on the function $d:X\times X\to\mathbb{R}$ to be a metric and instead allow any symmetric, measurable and bounded function $k\in L^{p}_{\mathrm{sym}}(X\times X)$. Then, the structure $M=(X,k,\mu)$ is instead referred to as a gauged measure space [43] or a measure network [10] (in fact, the definition of a measure network in [10] even drops the symmetry condition). This allows us to consider pairwise relations on $X$ to be modelled by a wider class of functions, such as affinity kernels. In this case, the space $(\mathcal{M},d_{\mathrm{GW},p})$ of equivalence classes of measure networks (under the equivalence relation $M\sim M^{\prime}\Leftrightarrow d_{\mathrm{GW},p}(M,M^{\prime})=0$) is a complete, geodesic, metric space [43, 10].
Recently, a GW-like distance has been developed for comparing hypergraph structures [11], based on ideas from the co-optimal transport problem introduced in [45]. Recall that a hypergraph is defined by a pair $H=(V,E)$, where $V$ is a finite set of vertices and $E$ is a set of hyperedges; each $e\in E$ is a subset of $V$. A general structure which encompasses the notion of a hypergraph is a measure hypernetwork, which is a quintuple $H=(X,\mu,Y,\nu,\omega)$, where $(X,\mu)$ and $(Y,\nu)$ are respectively Polish spaces with fully supported Borel probability measures, and $\omega$ is a non-negative, measurable and bounded function $\omega:X\times Y\to\mathbb{R}$. Indeed, from a hypergraph $(V,E)$, one obtains a measure network $(V,\mu,E,\nu,\omega)$, where $\mu$ and $\nu$ are uniform probability measures and $\omega:V\times E\to\mathbb{R}$ is a binary incidence function ($\omega(v,e)=1\Leftrightarrow v\in e$).
Given two measure hypernetworks $H=(X,\mu,Y,\nu,\omega)$ and $H^{\prime}=(X^{\prime},\mu^{\prime},Y^{\prime},\nu^{\prime},\omega^{\prime})$, the $p$-co-optimal distortion for a pair of couplings $(\pi^{v},\pi^{e})\in\Pi(\mu,\mu^{\prime})\times\Pi(\nu,\nu^{\prime})$ (the notation here is intended to evoke the idea that $\pi^{v}$ is a coupling of vertices and $\pi^{e}$ is a coupling of hyperedges) is defined to be
$\mathrm{dis}^{\mathrm{COOT}}_{p}(\pi^{v},\pi^{e}):=\left(\int_{X\times X^{\prime}\times Y\times Y^{\prime}}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{e}(y,y^{\prime})\right)^{1/p}=\|\omega-\omega^{\prime}\|_{L^{p}(\pi^{v}\otimes\pi^{e})}.$
Similarly to the GW distance, the hypernetwork $p$-co-optimal transport distance is given by
@@
The expression $d_{\mathrm{COOT},p}(H,H^{\prime})$ defines a pseudo-metric on the space of measure hypernetworks, and induces a metric on the space of equivalence classes of measure hypernetworks up to isomorphism, which is also shown to be geodesic and complete in [11, Theorem 1].
@@ 3 Topological Optimal Transport
Having discussed various matching-based comparison approaches for persistence diagrams, point clouds, and hypergraphs, we turn to the aim of this article, which is to develop an optimal transportation framework for matching point clouds that couples geometric and topological information. Our starting point is to represent a point cloud $X$ together with its topological features as a PH-hypergraph $H=H(X,\mathcal{K},g)$. Given another point cloud and PH-hypergraph, our objective is to find a coupling between them that minimises topological distortion and optimally preserves topological features. A natural attempt to a find a topology-informed matching between point clouds $X$ and $X^{\prime}$ is to simply apply the HyperCOT framework [11] on their PH-hypergraphs $H=H(X,\mathcal{K},g)$ and $H^{\prime}=H^{\prime}(X^{\prime},\mathcal{K}^{\prime},g^{\prime})$. However, this simplistic approach does not work due to two problems we list below.
  • Problem 1. PH-hypergraphs do not contain any information on the significance (measured by persistence) of homology classes. One could naturally weight hyperedges by the persistence value of their corresponding homology classes [2]. However, in an optimal transport setting, the HyperCOT would promote “splitting” of mass from hyperedges with higher persistence to hyperedges with lower persistence, rather than matching significant hyperedges.
  • Problem 2. PH-hypergraphs can be disconnected, and there might be points which do not belong to any hyperedge [2]. In this context, the desired property is to have geometric (spatial) information co-driving the transport plan.
@@ 3.1 Measure topological networks
To solve the problems described above, we first introduce an appropriate general model for the structures that we wish to compare. Here we introduce the following notation. For a locally compact Polish metric space $Y$, we denote by $\overline{Y}$ the augmented space obtained as the disjoint union $\overline{Y}=Y\cup\{\partial_{Y}\}$ (with the disjoint union topology), where $\partial_{Y}$ is an abstract point.
@@
Let us denote by $\mathcal{P}$ the class of all measure topological networks, with the simplifying conventions that gauged measure spaces have bounded kernels and that measure persistence diagrams have finite $\infty$-persistence.
@@
A measure topological network arises from the data of a PH-hypergraph defined over a metric measure space, and this is inspiration for the definition and the source of all computational examples. Indeed, given a finite metric measure space $(X,d,\mu)$, an associated persistence diagram $D$ (obtained via, say, Vietoris-Rips persistent homology) and a choice of a set of generating cycles $g$ for $D$, we construct an associated topological network as follows. First, we take $Y=D$ as a multiset of points in $\Lambda$ (one can consider this as a proper set by indexing its points, so that $Y$ can be considered a finite topological space). Taking $\nu$ to be the uniform (i.e., counting) measure and $\iota:Y\to\Lambda$ to be inclusion, we have that $\iota_{\#}\nu=\nu_{D}$ as in (3). Finally, $\omega$ is a kernel representing the hypergraph structure $(X,g)$. In our computational experiments, we use the binary incidence function: for $x\in X$ and a point $(a,b)\in D$, represented by a cycle $c\in g$, $\omega(x,c)=1$ if and only if $x$ is a vertex of any simplex in $c$. The flexibility in the definition allows for other hypergraph kernels, such as those studied in [11]. We can also replace the metric $d$ with a symmetric gauge function $k$ that may better encode local geometry without issue—see Section 5 for examples.
@@
In full generality, topological measure networks are not restricted to describing topological features of the underlying space. The construction described in Example 1 motivates the use of “topology” in the name.
@@
We now define a notion of distance between measure topological networks. This distance will be our main object of study throughout the rest of the paper.
@@
$\qquad\qquad\qquad\qquad+\int_{(X\times X^{\prime})^{2}}|k(x,y)-k^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{v}(y,y^{\prime})$ (7)
$\qquad\qquad\qquad\qquad\left.+\int_{X\times X^{\prime}\times\overline{Y}\times\overline{Y^{\prime}}}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{e}(y,y^{\prime})\right)^{\frac{1}{p}}.$ (8)
@@
$d_{\mathrm{TpOT},p}(P,P^{\prime})=\inf_{\begin{subarray}{c}\pi^{v}\in\Pi(\mu,\mu^{\prime})\\ \pi^{e}\in\Pi_{\rm adm}(\nu,\nu^{\prime})\end{subarray}}\left(\|d_{p}\circ(\iota,\iota^{\prime})\ \|^{p}_{L^{p}(\pi^{e})}+\|k-k^{\prime}\|^{p}_{L^{p}(\pi^{v}\otimes\pi^{v})}+\|\omega-\omega^{\prime}\|^{p}_{L^{p}(\pi^{v}\otimes\pi^{e})}\right)^{\frac{1}{p}}.$ (9)
Similar to Section 2.2, we say a Radon measure $\pi$ on $\overline{Y}\times\overline{Y^{\prime}}$ is admissible, and we write $\pi\in\Pi_{\mathrm{adm}}(\nu,\nu^{\prime})$, if it has marginals $\nu$ and $\nu^{\prime}$, and it satisfies $\pi((\partial_{Y},\partial_{Y^{\prime}}))=0$. Here we are implicitly extending $\omega$ and $\iota$ to functions $\omega:X\times\overline{Y}\to\mathbb{R}$, $\iota:\overline{Y}\to\overline{\Lambda}$ by setting $\omega_{|X\times\partial_{Y}}=0$ and $\iota(\partial_{Y})=\partial_{\Lambda}$. The restriction $\omega_{|X\times\partial_{Y}}$ can be interpreted as (trivial) a membership function (via the map $\iota$) for the diagonal $\partial_{\Lambda}$.
@@
  • The integral in (6) is a Wasserstein term which measures the quality of the matching of generating cycles. This term is intended to address Problem 1.
  • The integral in (7) is a Gromov-Wasserstein term which matches the quality of the matching of points in the gauged measure spaces. This term is intended to address Problem 2.
  • The integral in (8) is a co-optimal transport term which addresses the basic problem of matching PH-hypergraphs.

### v2

@@ 2.2 Optimal transport on persistence diagrams
where the $\delta_{x_{i}}$ denotes a Dirac supported at $x_{i}=(b_{i},d_{i})$.
Let $\overline{\Lambda}:=\Lambda\cup\{\partial_{\Lambda}\}$ be obtained by adjoining to $\Lambda$ the virtual point $\partial_{\Lambda}$, with the resulting space endowed with the disjoint union topology. Given two measure persistence diagrams $\nu$ and $\nu^{\prime}$, we say a Radon measure $\pi$ on $\overline{\Lambda}\times\overline{\Lambda}$ is admissible if it has marginals $\nu$ and $\nu^{\prime}$ (we consider $\nu$ and $\nu^{\prime}$ to be measures on $\overline{\Lambda}$ which are supported on $\Lambda$) and it satisfies the additional constraint $\pi((\partial_{\Lambda},\partial_{\Lambda}))=0$, and write $\pi\in\Pi_{\mathrm{adm}}(\nu,\nu^{\prime})$. Then for $p\in[1,\infty)$, we define the $p$-Wasserstein distance between measure persistence diagrams [19] as:
$d_{\mathrm{W},p}^{\mathrm{MPD}}(\nu,\nu^{\prime})=\inf_{\pi\in\Pi_{\rm adm}(\nu,\nu^{\prime})}\left(\int_{\overline{\Lambda}\times\overline{\Lambda}}\|x-x^{\prime}\|^{p}_{p}\>\mathrm{d}\pi(x,x^{\prime})\right)^{1/p}.$ (4)
If $\nu_{D}$ and $\nu_{D^{\prime}}$ arise from finite persistence diagrams $D$ and $D^{\prime}$, respectively, as in (3), then it is not hard to show that $d_{\mathrm{W},p}^{\mathrm{PD}}(D,D^{\prime})=d_{\mathrm{W},p}^{\mathrm{MPD}}(\nu_{D},\nu_{D^{\prime}})$ [19, Proposition 3.2]. In general, $d_{\mathrm{W},p}^{\mathrm{MPD}}$ is finite, provided that we work in the space of measure persistence diagrams $\mu$ with finite $p$-persistence, in the sense that
$\left(\int_{\Lambda}\|x-\mathrm{Proj}_{\partial_{\Lambda}}(x)\|^{p}_{p}\>\mathrm{d}\mu(x)\right)^{1/p}<\infty.$ (5)
In what follows, we will restrict to the space of measure persistence diagrams $\mu$ with finite $p$-persistence for every $p\geq 1$. This is mostly for convenience, to avoid the need to include additional technical conditions and notation in the statements of our results. With an abuse of notation, we will still indicate the space of such objects as $\mathrm{MPD}$.
@@ 2.3 Gromov-Wasserstein and co-optimal transport distances
A metric measure space (mm-space) [54, 42] is a triple $M=(X,d,\mu)$, where $X$ is a space endowed with a complete separable metric $d$ and with a fully supported Borel measure $\mu$. Given two mm-spaces $M=(X,d,\mu)$ and $M^{\prime}=(X^{\prime},d^{\prime},\mu^{\prime})$ and a coupling $\pi\in\Pi(\mu,\mu^{\prime})$ and $p\in[1,\infty)$, we introduce the $p$-distortion functional $\mathrm{dis}_{p}^{\mathrm{GW}}$ [54]:
@@
For our purposes, we relax the requirement on the function $d:X\times X\to\mathbb{R}$ to be a metric and instead allow any symmetric, measurable and bounded function $k\in L^{p}_{\mathrm{sym}}(X\times X)$. Then the structure $M=(X,k,\mu)$ is instead referred to as a gauged measure space [54] or a measure network [11] (in fact, the definition of a measure network in [11] even drops the symmetry condition). This allows us to consider pairwise relations on $X$ to be modelled by a wider class of functions, such as affinity functions or kernels [12, 65]. In this case, the space $(\mathcal{M},d_{\mathrm{GW},p})$ of equivalence classes of measure networks (under the equivalence relation $M\sim M^{\prime}\Leftrightarrow d_{\mathrm{GW},p}(M,M^{\prime})=0$) is a complete, geodesic, metric space [54, 11].
Recently, a GW-like distance has been developed for comparing hypergraph structures [13], based on ideas from the co-optimal transport problem initially introduced in [59]. Recall that a hypergraph is defined by a pair $H=(V,E)$, where $V$ is a finite set of vertices and $E$ is a set of hyperedges; each $e\in E$ is a subset of $V$. A general structure which encompasses the notion of a hypergraph is a measure hypernetwork, which is a quintuple $H=(X,\mu,Y,\nu,\omega)$, where $(X,\mu)$ and $(Y,\nu)$ are respectively Polish spaces with fully supported Borel probability measures, and $\omega$ is a non-negative, measurable and bounded function $\omega:X\times Y\to\mathbb{R}$. Indeed, any hypergraph $(V,E)$ can be encoded as a measure network $(V,\mu,E,\nu,\omega)$, by choosing $\mu$ and $\nu$ to be uniform probability measures on $V$ and $E$ respectively, and $\omega:V\times E\to\mathbb{R}$ to be an indicator $\omega(v,e)=\mathbf{1}_{v\in e}$.
Given two measure hypernetworks $H=(X,\mu,Y,\nu,\omega)$ and $H^{\prime}=(X^{\prime},\mu^{\prime},Y^{\prime},\nu^{\prime},\omega^{\prime})$, the $p$-co-optimal distortion for a pair of couplings $(\pi^{v},\pi^{e})\in\Pi(\mu,\mu^{\prime})\times\Pi(\nu,\nu^{\prime})$ (the notation here intended to evoke the idea that $\pi^{v}$ is a coupling of vertices and $\pi^{e}$ is a coupling of hyperedges) is defined to be
$\mathrm{dis}^{\mathrm{COOT}}_{p}(\pi^{v},\pi^{e})$ $:=\left(\int_{X\times X^{\prime}\times Y\times Y^{\prime}}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{e}(y,y^{\prime})\right)^{1/p}$
$\>=\|\omega-\omega^{\prime}\|_{L^{p}(\pi^{v}\otimes\pi^{e})}.$
Similarly to the GW distance, the hypernetwork $p$-co-optimal transport distance is then given by
@@
The expression $d_{\mathrm{COOT},p}(H,H^{\prime})$ defines a pseudo-metric on the space of measure hypernetworks, and induces a metric on the space of equivalence classes of measure hypernetworks up to isomorphism, which is also shown to be geodesic and complete in [13, Theorem 1].
@@ 3 Topological Optimal Transport
Having discussed matching-based comparison approaches for persistence diagrams, point clouds, and hypergraphs respectively, we turn to the aim of this article, which is to develop a framework for matching point clouds that couples geometric and topological information via these structures. Our starting point is to represent a point cloud $X$ together with its topological features as a PH-hypergraph $H=H(X,\mathcal{K},g)$. Given another point cloud and PH-hypergraph, our objective is to find a coupling between them that minimises topological distortion and optimally preserves topological features, in some sense. A naïve attempt to a find a topology-informed matching between point clouds $X$ and $X^{\prime}$ is to simply apply the HyperCOT framework [13] on their PH-hypergraphs $H=H(X,\mathcal{K},g)$ and $H^{\prime}=H^{\prime}(X^{\prime},\mathcal{K}^{\prime},g^{\prime})$. However, this simplistic approach suffers from at least two problems which we list below.
  • Problem 1. PH-hypergraphs do not contain any information on the significance (measured by persistence) of homology classes. One could naturally weight hyperedges by the persistence value of their corresponding homology classes [2]. However, in an optimal transport setting, the HyperCOT would promote “splitting” of mass from hyperedges with higher persistence to hyperedges with lower persistence, rather than the desired effect of matching significant hyperedges.
  • Problem 2. PH-hypergraphs can be disconnected, and there might be points which do not belong to any hyperedge [2]. In this context, the desired property is to have geometric (spatial) information inform the transport plan.
@@ 3.1 Measure topological networks
To solve the problems described above, we first introduce an appropriate general model for the structures that we wish to compare. We first introduce necessary notation as follows: for a locally compact Polish metric space $Y$, we denote by $\overline{Y}$ the augmented space obtained as the disjoint union $\overline{Y}=Y\cup\{\partial_{Y}\}$ (with the disjoint union topology), where $\partial_{Y}$ is an abstract point.
@@
Let us denote by $\mathcal{P}$ the class of all measure topological networks, with the simplifying conventions that gauged measure spaces have bounded kernels and that measure persistence diagrams have finite $p$-persistence for all $p\geq 1$ (see (5)).
@@
A measure topological network arises from the data of a PH-hypergraph defined over a metric measure space, and this is inspiration for the definition and the source of all computational examples. Indeed, given a finite metric measure space $(X,d,\mu)$, an associated persistence diagram $D$ (obtained via, say, Vietoris-Rips persistent homology) and a choice of a set of generating cycles $g$ for $D$, we construct an associated topological network as follows. First, we take $Y=D$ as a multiset of points in $\Lambda$ (one can consider this as a proper set by indexing its points, so that $Y$ can be considered a finite topological space). Taking $\nu$ to be the uniform (i.e., counting) measure and $\iota:Y\to\Lambda$ to be inclusion, we have that $\iota_{\#}\nu=\nu_{D}$ as in (3). Finally, $\omega$ is a kernel representing the hypergraph structure $(X,g)$.
The simplest choice is to use a binary incidence function: for $x\in X$ and a point $(a,b)\in D$, represented by a cycle $c\in g$, $\omega(x,c)=1$ if and only if $x$ is a vertex of any simplex in $c$. The flexibility in the definition allows for other hypergraph kernels which may have more desirable properties in practice, such as those studied in [13]. Of note, the recently developed method TOPF [31] produces node-level representations of topological features in the form of a scalar function on points and topological features, and as such fits within our framework. Finally, we can replace the metric $d$ with a symmetric gauge function $k$ that may better encode local geometry. We refer the reader to Section 5 for examples of different choices of incidence and gauge functions.
@@
In full generality, measure topological networks are not restricted to describing topological features of the underlying space. The construction described in Example 1 motivates the use of “topological” in the name.
@@
Remark 2 (Finite and Infinite Diagrams).
We acknowledge that our main examples of topological networks from Example 1 (as well as the examples described in Example 2) have the property that $Y$ is finite; that is, the relevant MPDs are finitely supported. One could safely work under this assumption throughout the analysis in the rest of the paper. However, this assumption doesn’t make any of the analysis significantly easier, so we opt to work at the level of generality which allows for sets $Y$ with infinite cardinality. Moreover, there are interesting examples where non-finitely supported MPDs arise, and we prefer to work with a formalism which is able to handle these examples, with a view toward future applications. One source of such examples is the statistical theory of the topology of random point clouds, where non-finite MPDs play an important role as asymptotic distributions when the number of points tends to infinity [5, 20, 28].
We now aim to define a notion of distance between measure topological networks, which will be the main object of study throughout the rest of the paper. First, we synthetically extend the definition of an admissible coupling from the setting of persistence diagrams (see Section 2.2) to the general setting of measure spaces with adjoined virtual points.
Definition 2 (Admissible Couplings).
Let $(Y,\nu)$ and $(Y^{\prime},\nu^{\prime})$ be a locally compact Polish spaces endowed with a Radon measures and let $\overline{Y}=Y\cup\{\partial_{Y}\}$ and $\overline{Y^{\prime}}=Y^{\prime}\cup\{\partial_{Y^{\prime}}\}$ be the associated augmented spaces. We say that a Radon measure $\pi$ on $\overline{Y}\times\overline{Y^{\prime}}$ is admissible if it has marginals $\nu$ and $\nu^{\prime}$ and satisfies $\pi((\partial_{Y},\partial_{Y^{\prime}}))=0$. We denote the set of admissible couplings by $\Pi_{\mathrm{adm}}(\nu,\nu^{\prime})$.
This leads to the following main definition.
@@
$\quad\quad\quad+\int_{(X\times X^{\prime})^{2}}|k(x,y)-k^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{v}(y,y^{\prime})$ (8)
$\quad\quad\quad\left.+\int_{X\times X^{\prime}\times\overline{Y}\times\overline{Y^{\prime}}}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{e}(y,y^{\prime})\right)^{\frac{1}{p}}.$ (9)
@@
$\begin{split}d_{\mathrm{TpOT},p}(P,P^{\prime})&=\inf_{\pi^{v},\pi^{e}}\left(\|d_{p}\circ(\iota,\iota^{\prime})\ \|^{p}_{L^{p}(\pi^{e})}\right.\\ &\qquad\qquad\qquad+\left.\|k-k^{\prime}\|^{p}_{L^{p}(\pi^{v}\otimes\pi^{v})}+\|\omega-\omega^{\prime}\|^{p}_{L^{p}(\pi^{v}\otimes\pi^{e})}\right)^{\frac{1}{p}}.\end{split}$ (10)
Here we are implicitly extending $\omega$ and $\iota$ to functions $\omega:X\times\overline{Y}\to\mathbb{R}$, $\iota:\overline{Y}\to\overline{\Lambda}$ by setting $\omega\lvert_{X\times\partial_{Y}}=0$ and $\iota(\partial_{Y})=\partial_{\Lambda}$. The restriction $\omega\lvert_{X\times\partial_{Y}}$ can be interpreted as (trivial) a membership function (via the map $\iota$) for the diagonal $\partial_{\Lambda}$.
@@
  • The integral in (7) is a partial transport term which measures the quality of the matching of generating cycles. This term is intended to address Problem 1.
  • The integral in (8) is a “Gromov-Wasserstein” term which matches the quality of the matching of points in the gauged measure spaces. This term is intended to address Problem 2.
  • The integral in (9) is a “co-optimal transport” term which addresses the basic problem of matching PH-hypergraphs.